OpenStudy (anonymous):

How do you find the distance between the following Cartesian Plane coordinates; (2, -7) and (6, 1)? Give your answer in exact simplest form.

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

But how do I know what is x1 and which is x2 etc?

hartnn (hartnn):

2,-7 x1 = 2 y1 = -7

OpenStudy (anonymous):

Where \[(x_1,y_1)=(2,-7)\] \[(x_2,y_2)=(6,1)\]

hartnn (hartnn):

6,1 x2 = 6 y2 = 1

OpenStudy (anonymous):

Does it matter which order the numbers are placed in?

hartnn (hartnn):

only if its same co-ordinate like x1 - x2 or x2 -x1 , you will get same result similarly, you can do, y2-y1 or y1-y2

hartnn (hartnn):

thats because you have to square them like 4-2 = 2 squaring gives 4 2-4 = -2 squaring again gives 4

OpenStudy (anonymous):

you can choose any pair of co-ordinates as x1,y1 then the other is x2,y2 .

OpenStudy (whpalmer4):

This is just the Pythagorean theorem in disguise:  You're just interested in the magnitude of the side lengths of the triangle. You could write the formula using absolute value notation, but because you square the side lengths, you always get a positive number and it isn't necessary to write the absolute value symbols.